You are here: Home / Documents / Minutes of CASU Meetings / minutes_100630.txt

minutes_100630.txt

Plain Text icon minutes_100630.txt — Plain Text, 11 kB (11548 bytes)

File contents

Wednesday 30th June 2010   APM Meeting tables    14:00 - 15:00

Present:     MJI, EGS, JRL, RGM, SC, JCR
Apologies:   NAW, MR, STH, JPE


Agenda
=======

1.  Actions from last meeting
2.  Comments on WFAU minutes
3.  Recent & upcoming meets
4.  Data archives update     - AAT, ING, WFCAM, UKIRT, VISTA
5.  Optical/NIR processing   - HAWKI, INT WFC, MegaCam, Subaru, VST
6.  WFCAM update
7.  VISTA update
8.  AOB


Minutes
========

1. Actions from the last meeting
--------------------------------

EGS  stage I software release almost done just a bit of building left      <<<<
JRL 

STH  finishing off last bits of calibration plan document update           <<<<

RGM  still testing and noted a 1% difference at the 5-sigma level between 
     the different ways people are estimating the magnitude error due to
     most other folk doing it wrong                                        <<<<

ALL  contribute text to go with CASU versions of VISTA PR images - ongoing <<<<
     
EGS  contacted VVV team to get an example of their multi-step tiles for 
     comparison.  On visual inspection they look very similar but their
     tile astrometry is a little odd

JPE  succeeded!? in influencing the powers that be on the IOT duration

JPE  contact ESO re: signing off pipeline deliverables - ongoing           <<<<


2. WFAU minutes
----------------

Not many but JRL remembered he'd heard nothing back about the tile products
yet.  MJI said that he would reply about the photometric corrections 
viz-a-viz tiles mentioned in the latest minutes.                           <<<<

MJI asked RGM to fill us in on the recent PSPI telecon organised by NJC.
As far as he could recall there were no CASU issues that were raised.
MJI has a copy of the minutes if anyone is interested.


3. Recent and Upcoming Meetings
-------------------------------

VISTA IOT telecon: new & improved at a mere 1.25 hours.  MJI sat in on it
as did JPE in San Diego.  Since the previous intervention there has been a 
problem with elliptical images.  ESO undertook a further intervention 
between 16-21st June to address this and a series of other issues. We now 
apparently have round images again but possibly still a small focus gradient 
affecting some detectors.  He also noted that the calibration units have
been installed and after a few flux level tests good linearity sequences 
et al. should apparently be winging our way.  This together with the news
that a set of flats for June have been taken recently means that we can
begin processing May/June data soon.

There was a brief discussion at the telecon about OB classification for
standards which if MJI's memory serves seems to be X.  MJI also learnt
that Wolfgang had compiled a report on persistence which he asked to
get a copy of.  The consensus seemed to be that, like other NIR arrays,
the persistence was not quantitatively predictable and that correcting
for it was therefore impractical.  Since, as noted before, a sensible
observing sequence precludes it being a problem ...... its not a problem.

MJI also learnt that the sequence of missing darks for VISTA in June was 
caused by the script that sorts this going walkabout after some recent 
system changes.

MJI duly trained down to Imperial and presented a talk at the UKIRT board 
meeting on 10th June.  This seemed to be well received.

MJI noted the upcoming: "UKIDSS and VISTA Galactic surveys meeting",
organised by Melvin Hoare, in Leeds on Thursday 8th July 11am-6pm.  JPE and 
MJI are going but if anyone else fancies a day out in Leeds, you know where
it is.

There is an EGAPS meeting in Cambridge on 12th July, followed on the 13th
by a more detailed review of progress for the IPHAS global calibration 
starring Janet Drew and Brent Miszalski.

Next week, on **Tuesday 6th July**, we will have an internal half-day
review (9am-1pm with a long coffee break) of VISTA pipe-line processing 
from raw data to calibrated science products for our new and recent recruits, 
and all other interested local parties (with only a modest entrance fee).  <<<<

NAW sent in the following report on money-laden (i.e. space-based activities):

PLATO - the ESA AO will be issued in July - thus there will be some 
work involved in providing input into this in terms of the development 
and planning for the Exoplanet processing system - where the IoA/CASU 
has lead responsibility

Gaia - development of the case Post Launch Support grant is now 
starting up - this is expected to be required by STFC/UKSpA for 
evaluation at the end of 2010.

GREAT - plenary meeting #3 development of the GCDS survey - 
identified the need for input photometry for catalogue generation requiring
access to mainly CASU generated WFCAM/generic optical and VISTA/VST data.
 
MJI for CASU, Colin Vincent for STFC and Bob Mann and Keith Noddle for WFAU
duly turned up at ESO on 2nd June to attempt to further discussion regarding
ESO in-kind repayment possibilities with Michele, Fernando, Pascal, Andreas
and Martino.  The ESO council meeting subsequent to this dead-batted the ball
back down the wicket so ...... once more unto the breach dear friends.


4. Data Archive
---------------

ING, AAO, UKIRT, VISTA  all up to date and trucking ....... well they were
until the brutish SSH attacks and retaliatory measures taken last week.  
A sweepstake was organised on who will complain first.


5. Optical / NIR Processing
---------------------------

Nothing mentioned nothing gained so presumably all chugging along fine.  
A bit of IPHAS'ing re: global calibration and a bit of Subaru processing
from EGS otherwise all quiet once more in the Hobyar House.


6. WFCAM Update
---------------

WFCAM was taken off for a well-earned rest for CASU from the 25th May to 9th 
June.  Oddly enough the UKIRT primary mirror just happened to be re-aluminised
at the same time.  

Raw data ingest and verification is up to date but WFCAM processing has been 
on hold since the re-aluminising awaiting a full set of twilight flats.  The 
K-band flats had to be retaken on the 29th June but after that processing was 
fired up full steam and should soon be back on track.  One night (28th) has a 
dropped out camara #3 for the whole night and since all the MSBs were 
rejected we will flag it as unprocessed for now.

There has been a further shrinkage of staff at UKIRT.  Tom Kerr is now our
main contact with Watson the other UKIRT contact.  Chris has moved on to
JCMT matters.

MJI heard from Will (Hartely) and Omar that the reprocessed 05b-07b J-band
UDS data looks a lot better than the original so as soon as feasible it 
looks like we will begin redoing the equivalent H-band set.


7. VISTA Update
---------------

Ingestion and verification of data up to 19th June is in place but we have 
no suitable flats or linearity sequences yet to begin processing after
the intervention at the end of April.  JRL noted that the reprocessing of 
data (v1.0 with tiles) up to 23rd April will be completed by the end of this
week, honest !

MJI asked EGS if he could add on the April Zpt info to the current Zpt
monitoring plot, complete up to end of March to see if the trends of Zpt with
time are flattening off or reversing due to seasonal effects.              <<<<

NaNs et al.: WFAU found a minor problem with a catalogue that JRL is 
investigating - 10 stars had values of 0 across the board, including
object ID.  Can't possibly happen = needs some investigating.

RGM noted that how, and particularly when, data will be delivered to ESO
is still not time-tabled (this was discussed at the PSPI telecon noted
earlier). There is to be a Phase III viz 'How to deliver data to ESO' type 
workshop at some point soon(ish) allegedly.

MJI then asked for a compilation of currently known FEATURES in the v1.0
science products some of which need to be fixed prior to releasing them.
As these are all minor and involve trivial header faux pas this should
be straightforward:

* nice to see the EXTNAMES in VIRCAM products are now there but in the 
  pawprint catalogues they are now duplicated with both APM-BINARYTABLE and 
  DET1.CHIPn etc...The image products are fine.                            <<<<
* some required PHU header info in the tile catalogue SHUs is missing
  this is needed for compatibility with the pawprint catalogues            <<<<
* the QC OB info still needs inserting into the FITS headers               <<<<
* the astrometric ZPN constants are slightly different from the latest values
  derived but until we have seen what the effect of the recent interventions
  has been there is no point in updating them yet
* likewise the photometric colour equations may need a slight tweak which
  awaits STH's further investigations

MJI noted that we need to revisit the SV data encore une fois after v1.0 
dust has settled.                                                          <<<<


8. AOB
-------

Due to the ongoing brute force attack on SSH there was a discussion
about how best to make sure that those who need to can still log on
and run the pipelines etc... remotely. At present some APM computers are
accessible but not others. SC will look into this. Meanwhile, using
VPN through the University system is being investigated.                   <<<<

An eye needs to be kept on disk space for apm29/35 since they have filled
up twice now with all this reprocessing kerfuffle.

It's hot (JCR) pleasant (MJI) down here - upper windows can be kept open 
24/7 to keep the punters happier.

New disk arrays are going to be needed for VISTA soon and we need to decide 
whether to get the same as usual or investigate other options.

The apm 29 & 35 disk arrays seem to be running standard book-keeping jobs 
painfully slowly (e.g. rm's, chmod -R's) but otherwise they are fine. This
requires further investigation, is it a feature of ZFS ???                 <<<<

Only DWE whinged about the Plone upgrade complaining it crashed his
latest white hot flame of technology Safari browser.  There's a moral there,
just not obvious what it is.

MJI suggested that he, NAW and SC get together to discuss the management
of the non-astro machines.                                                 <<<<

Finally MJI asked SC if she could order some more tapes                    <<<<


Continuing Actions:
===================

EGS  - finish adding catalogue software to release pages
JRL

STH  - finish off last bits of calibration plan document update
JRL

RGM  - finish off testing catalogue conversion software

ALL  - contribute text to go with CASU versions of VISTA PR images
       add links to ESO versions

JPE  - contact ESO re: signing off pipeline deliverables


New Actions:
============

MJI  - contact WFAU about tile photometric features

ALL  - prepare/revamp some viewgraphs for the review on 6th July

EGS  - update Zpt trend plot on technical web page

JRL^3- Jim'll fix FEATURES listed earlier

EGS  - SV time again
STH
MJI

SC   - investigate University VPN route and check all apm machines have 
MJI    limited access and that we can still access them remotely as needed

SC   - does our ZFS need a good tune ?

MJI  - talk to NAW and Asif about management of the the non-astro machines
SC

SC   -  order yet more LTO-II and IV tapes and longer ethernet cables